OpenStudy (anonymous):

Hillary jogged a total distance of 7 and 1 over 3 miles during the months of October and November. If Hillary only jogged 1 over 6 mile every day, which expression shows the number of days in which she went jogging?

OpenStudy (mathstudent55):

Where are the expressions?

OpenStudy (anonymous):

A. 7 1/3 * 1/6 B. 7 1/3 + 1/6 C. 7 1/3 - 1/6 D. 7 1/3 / 1/6

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@EclipsedStar

jimthompson5910 (jim_thompson5910):

\[\large \text{Total Distance} = (\text{miles jogged per day})*(\text{number of days})\] Divide both sides by "miles jogged per day" to get \[\large \text{Total Distance} = (\text{miles jogged per day})*(\text{number of days})\] \[\large \frac{\text{Total Distance}}{\text{miles jogged per day}} = \frac{(\text{miles jogged per day})*(\text{number of days})}{\text{miles jogged per day}}\] \[\large \frac{\text{Total Distance}}{\text{miles jogged per day}} = \text{number of days}\] \[\large \text{number of days} = \frac{\text{Total Distance}}{\text{miles jogged per day}}\]